Describe how to translate the graph y=sqrt(x) to obtain the graph of y=sqrt(x+5)
Please help with this math problem. I will appreciate your kind response.

Found 2 solutions by stanbon, graphmatics:
Answer by stanbon(75887)

x=sqrt(3x+40)
---
Square both sides to get:
x^2 = 3x+40
Rearrange:
x^2 - 3x - 40 = 0
---
x^2-8x+5x-40 = 0
x(x-8)+5(x-8) = 0
(x-8)(x+5) = 0
--------------
x = 8 or x = -5
---
Checking:
8 = sqrt(24+40)
8 = sqrt(64)
true
---
-5 = sqrt(-15+40)
false
====================
Only solution: x = 8
=========================
Cheers,
Stan H.

Answer by graphmatics(170)

Since they are both graphs of square root expression we should a a value to one expression to translate that expression to the other

sqrt(x)+z = sqrt(x+5)
z = sqrt(x+5)-sqrt(x)
So
to translate y = sqrt(x) simply add sqrt(x+5)-sqrt(x)
to get
y=sqrt(x+5)